OpenStudy (elenathehomeschooler):

Someone help! Write a function with the given characteristics. A polynomial with rational coefficients having roots 4 and 3-i

OpenStudy (elenathehomeschooler):

@jim_thompson5910 can you help?

jimthompson5910 (jim_thompson5910):

Hopefully you agree that if 3-i is one root, another root is 3+i, right?

OpenStudy (elenathehomeschooler):

yes i agree wch would make our roots 4 , 3-i , and 3+i right?

OpenStudy (elenathehomeschooler):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

if 4 is a root, then do you agree that x-4 is a factor?

OpenStudy (elenathehomeschooler):

yes

jimthompson5910 (jim_thompson5910):

if 3+i and 3-i is a root, then what must the factor be? any idea?

OpenStudy (elenathehomeschooler):

3?

jimthompson5910 (jim_thompson5910):

If 3+i is a root, then... x = 3+i x - 3 = i (x - 3)^2 = i^2 (x - 3)^2 = -1 (x - 3)^2 + 1 = 0 Hopefully you see how if 3+i is a root, then (x - 3)^2 + 1 is factor

OpenStudy (elenathehomeschooler):

ok

jimthompson5910 (jim_thompson5910):

agreed? or no?

OpenStudy (elenathehomeschooler):

yes i agree

jimthompson5910 (jim_thompson5910):

so we have these two factors (x-4) and ((x - 3)^2 + 1)

jimthompson5910 (jim_thompson5910):

(x - 3)^2 + 1 is the same as x^2-6x+9+1 = x^2-6x+10

OpenStudy (elenathehomeschooler):

would we end up FOIL them?

jimthompson5910 (jim_thompson5910):

so you need to multiply out (x-4) and (x^2-6x+10)

OpenStudy (elenathehomeschooler):

would it be x^3-10x^2+34x-40

jimthompson5910 (jim_thompson5910):

that is the correct answer

OpenStudy (elenathehomeschooler):

so that would be the final answer?

jimthompson5910 (jim_thompson5910):

yes